OpenStudy (anonymous):

simplify 3^1/2 x 3^1/2

OpenStudy (anonymous):

@skullpatrol

OpenStudy (solomonzelman):

\[a^{1/2} \times a^{1/2}\]when you multiply you add exponents so it will be \[a~^{(1/2~~~~+~~~~1/2)}=a^1=a\] Then, \[3^{1/2} \times 3^{1/2}=~?\]

OpenStudy (anonymous):

so it would be 3?

OpenStudy (anonymous):

i mean 9 lol

OpenStudy (solomonzelman):

No the first time you were correct it's 3.

OpenStudy (solomonzelman):

see why it's 3 ?

OpenStudy (anonymous):

really???

OpenStudy (anonymous):

hmmm no....

OpenStudy (anonymous):

i thought because its 3^1x3^1=9?

OpenStudy (solomonzelman):

But it's not ^1 it's ^(1/2)

OpenStudy (anonymous):

oh lol blonde moment XD

OpenStudy (solomonzelman):

\[\huge\color{red}{x^c \times x^d=}x^{(c+d)}\]

OpenStudy (anonymous):

thanks!

OpenStudy (solomonzelman):

in your case "c" and "d" are the same

OpenStudy (solomonzelman):

Anytime!